What would the concentration of acetic acid need to be in a solution containing 0.25 M...

  1. What would the concentration of acetic acid need to be in a solution containing 0.25 M sodium acetate to have a pH of 4.0? Given that Ka for acetic acid is 1.8 × 10-5.
Answer #1

Given:
Ka = 1.8*10^-5

Use:
pKa = -log Ka
= -log (1.8*10^-5)
= 4.745

Now use:
pH = pKa + log ([CH3COO-]/[CH3COOH])
4.0 = 4.745 + log(0.25 / [CH3COOH])
log(0.25 / [CH3COOH]) = -0.745
0.25 / [CH3COOH] = 10^(-0.745)
0.25 / [CH3COOH] = 0.180
[CH3COOH] = 1.39 M

Answer: 1.4 M